# appAppendix References

### v5

@@ G.1 Attack Setting
  • Attacker’s Access:
    After the fine-tuning phase, the attacker gains access to the fine-tuned model $p_{\theta}(\bm{x}\mid s)$. This access allows the attacker to perform the following actions:
    • Input Manipulation: The attacker can input any text prompt $s\in\mathcal{S}_{s}$ into the model to generate corresponding images:
@@
    • Data Extraction: By systematically querying the model with the same set of random strings $\mathcal{S}_{s}$, the attacker can generate multiple images $\{\bm{x}_{i,j}\}_{j=1}^{N_{G}}$ for each string $s_{i}$. The goal is to identify images that closely resemble the memorized training samples:
@@
  • Assumptions:
    • Knowledge of Surrogate Strings: The attacker is aware of the set of random strings $\mathcal{S}_{s}$ used during the victim’s fine-tuning process.
    • Model Query Capability: The attacker can perform a large number of queries to the fine-tuned model with each random string $s_{i}$ to generate sufficient samples for variance analysis.
    • No Access to Training Data: The attacker does not have direct access to the victim’s training dataset $\mathcal{D}_{train}$.
  • Attack Goals:
    • Data Reconstruction: To reconstruct or approximate the training images $\bm{x}_{i}$ by exploiting the model’s memorization of the random string-image associations.
    • Privacy Breach: To demonstrate the potential privacy risks associated with fine-tuned diffusion models by extracting sensitive or proprietary training data.
G.2 Extended SIDE Method
@@
The Extended SIDE method operates on the principle that random strings, when used as surrogate conditions during fine-tuning, become tightly coupled with specific training images. This follows the results of our theoretical understanding in Theorem 1. The attacker exploits this behavior to extract training data by performing the following steps:
@@ G.3 Algorithmic Implementation
G.4 Experiments
To verify the effectiveness of the Extended SIDE, we conduct experiments on the subset of the Laion-2B dataset, finetuning the Stable Diffusion v1.5 using the LORA[18] method. We use the following metrics to evaluate the attack results: mean SSCD (M-SSCD), AMS (mid-similarity) and LIPIPS. The results are shown in Tab. 4. We use the generated images of original prompts as the Baseline method. The extracted examples are shown in Figure 8. As the results show, our method could effectively extract training data from fine-tuned DPMs.
@@ G.4 Experiments
As presented in Tab. 4, the Extended SIDE method outperforms the baseline across all evaluation metrics:

### v6

@@ E.1 Attack Setting
  • Attacker’s Access: After the fine-tuning phase, the attacker gains access to the fine-tuned model $p_{\theta}(\bm{x}\mid s)$. Such access enables the attacker to perform Input Manipulation. The attacker can input any text prompt $s\in\mathcal{S}_{s}$ into the model to generate corresponding images:
@@
  • Data Extraction: By systematically querying the model with the same set of random strings $\mathcal{S}_{s}$, the attacker can generate multiple images $\{\bm{x}_{i,j}\}_{j=1}^{N_{G}}$ for each string $s_{i}$. Our goal is to identify images that closely resemble the memorized training samples:
@@
  • Assumptions: Knowledge of Surrogate Strings: The attacker is aware of the set of random strings $\mathcal{S}_{s}$ used during the victim’s fine-tuning process. Model Query Capability: The attacker can perform a large number of queries to the fine-tuned model with each random string $s_{i}$ to generate sufficient samples for variance analysis. No Access to Training Data: The attacker does not have direct access to the victim’s training dataset $\mathcal{D}_{train}$.
  • Attack Goals: Data Reconstruction: To reconstruct or approximate the training images $\bm{x}_{i}$ by exploiting the model’s memorization of the random string-image associations. Privacy Breach: To demonstrate the potential privacy risks associated with fine-tuned diffusion models by extracting sensitive or proprietary training data.
E.2 Extended SIDE Method
@@
The Extended SIDE method is based on the principle that random strings, when used as surrogate conditions during fine-tuning, become strongly associated with specific training images. This follows from the results of our theoretical understanding in Theorem 1. The attacker exploits this behavior to extract training data by performing the following steps:
@@ E.3 Algorithmic Implementation
E.4 Experiments
To verify the effectiveness of the Extended SIDE, we conduct experiments on the subset of the Laion-2B dataset, finetuning the Stable Diffusion v1.5 using the LORA[24] method. We use the following metrics to evaluate the attack results: mean SSCD (M-SSCD), AMS (mid-similarity) and LIPIPS. The results are shown in Table 5. We use the generated images of original prompts as the Baseline method. The extracted examples are shown in Figure 8.
@@ E.4 Experiments
As presented in Table 5, the Extended SIDE method outperforms the baseline across all evaluation metrics: